If you need .75 pounds of cheese to top a pizza, and you have 4 pounds of cheese at home, will you have enough to make 6 pizzas? Write an inequality where p = pizzas.
Thanks very much.

Answer by richwmiller(17219)

Sorry you will have to cut down on the cheese or make fewer pizzas
4/6=2/3=.667
6*2/3=4 pounds of cheese
4.5/6=3/4=.75
.75*6=4.5 pounds of cheese needed for 6 pizzas
.75*x=4
5 1/3 pizza so you can make 5 complete pizzas with 4 pounds of cheese


4<=.75p